OpenStudy (anonymous):

Solve the equation. 9=27x A.9 B.3 C.81 D.1/3

OpenStudy (anonymous):

How many times does 9 go into 27? 27 divided by 9. Then convert it into a fraction.

OpenStudy (anonymous):

Or...try each one. 9 = 27(9) 9 = 27(3) 9 = 27(81) 9 = 27(1/3)